OpenStudy (anonymous):

The length of a rectangle is 5m more than twice its width, and the area of the rectangle is 88m^2 . Find the dimensions of the rectangle

OpenStudy (anonymous):

Get two equations let breadth be 'b' and length be'l' 1=2b+5 (2b+5)*b=88 2b^2+5b-88=0 2b^2+16b-11b-88=0 2b(b+8)-11(b+8)=0 (b+8)(2b-11)=0 b=5.5m l=16

OpenStudy (anonymous):

whered u get 16b and 11b?

OpenStudy (anonymous):

it is another way of writing 5b i.e. 16-11=5b are you familiar with factorization cause thats what i did ????

OpenStudy (anonymous):

ahhhh gotcha sorry got lost